OpenStudy (anonymous):

B is generated from the formula below. B = roundeddown((H + 1) / 10)) The following table shows that when H = 150 that B = 15. Without knowing the formula in advance how do you use just the information in the table to determine the formula? 150 15 165 16 181 18 199 20 219 22 241 24 265 26 291 29 320 32 352 35 387 38 425 42 467 46 513 51 564 56 620 62 682 68 750 75 825 82 907 90

OpenStudy (anonymous):

Try plotting a graph and determine the formula from the graph.

OpenStudy (anonymous):

I am not aware of how to determine a formula from a graph.

OpenStudy (mathstudent55):

First, plot the points. Now look at the graph, if it looks linear, calculate the slope by choosing two points on the graph. Also, read the y-intercept from the plot. Then find the equation using y = mx + b, whre m is the slope and b is the y-intercept.

OpenStudy (anonymous):

X Y Slope 150 15 165 16 0.067 181 18 0.125 199 20 0.111 219 22 0.100 241 24 0.091 265 26 0.083 291 29 0.115 320 32 0.103 352 35 0.094 387 38 0.086 425 42 0.105 467 46 0.095 513 51 0.109 564 56 0.098 620 62 0.107 682 68 0.097 750 75 0.103 825 82 0.093 907 90 0.098 I realized there was an error in the information and changed it however neither table showed the slope as constant. That would mean that I am not able to use y=mx+b because it is not a line correct?

OpenStudy (anonymous):

I am not sure if this helps... X Y 0 0.1 1 0.2 2 0.3 3 0.4 4 0.5 5 0.6 6 0.7 7 0.8 8 0.9 9 1 10 1 11 1 12 1 13 1 14 1 15 1 16 1 17 1 18 1 19 2 21 2 23 2 25 2 27 2 29 3 32 3 35 3 38 3 41 4

OpenStudy (anonymous):

Well, the formula you provided sure isn't correct, not even if you round the numbers. (For example, (165 + 1) / 10 would be 16,6, rounded 17. Your table says B = 16. What would give you the exact B values for your given Hs would be dividing H by 10 and then "cutting off" the remainder. (There's 2 cases where it wouldn't work exactly: row 4, row 5. In these two cases it seems like the B values were rounded up because they were "very close" to another value e.g. 19,9 = 20 and 21,9 = 22. This is not the same as rounding up, the program probably recognized 19,9 to b 20 and 21,9 to be 22. This is a mistake that happens with a lot of badly programmed apps. It's got to do with the binary system). How did you obtain these values and what exactly do you want to do/describe with them?

OpenStudy (anonymous):

You are correct. My apologies. 16.6 would have to be rounded down to 16 before being added.

OpenStudy (anonymous):

The values are not important to me. The understanding of how to take values like this and generate a formula to predict their future values.

OpenStudy (anonymous):

Row 4 + 5 are correct. (199 + 1) / 10 = 20 and (219 + 1) / 10 = 22.